OpenStudy (gabbyalicorn):

A juice bar conducted a survey to determine which juice drink its customers prefer. The juice bar selected every 10th customer in 1 week, with 75 customers surveyed in all. The juice bar made a table of the results. What inferences can be drawn from this survey? Choose all answers that are correct. A. About 17% of the customers like Power Pear the most. B. About 21% of the customers like Wild Berry the most. C. Lemon Lift is the juice drink preferred by most customers. D. Tangy Apple is the juice drink preferred by the fewe

OpenStudy (gabbyalicorn):

D. Tangy Apple is the juice drink preferred by the fewest customers.

OpenStudy (gabbyalicorn):

@Lokid

OpenStudy (gabbyalicorn):

@Compassionate

OpenStudy (anonymous):

take each # and divide by 75

OpenStudy (gabbyalicorn):

1o 1 and 75?

OpenStudy (anonymous):

that should give the answer

OpenStudy (gabbyalicorn):

to 1- 1 and 75?

OpenStudy (anonymous):

25/75=.333333..

OpenStudy (anonymous):

that sorta thing

OpenStudy (gabbyalicorn):

:I

OpenStudy (anonymous):

17/75=.226666... 12/75=.16 21/75=.28

OpenStudy (gabbyalicorn):

should we round

OpenStudy (anonymous):

now make them % yes

OpenStudy (anonymous):

33% 27% 16% 28%

OpenStudy (anonymous):

so which do u think?

OpenStudy (gabbyalicorn):

A & B

OpenStudy (anonymous):

D

OpenStudy (gabbyalicorn):

just D? euh :/

OpenStudy (anonymous):

and

OpenStudy (gabbyalicorn):

...... :3

OpenStudy (anonymous):

C and D i think

OpenStudy (gabbyalicorn):

"think"

OpenStudy (anonymous):

C D C D C D

OpenStudy (anonymous):

very sure

OpenStudy (gabbyalicorn):

how sure r i=u

OpenStudy (anonymous):

very

OpenStudy (gabbyalicorn):

what percent?

OpenStudy (anonymous):

huh?

OpenStudy (gabbyalicorn):

1 to 100 how sure r u

OpenStudy (anonymous):

99.99%

OpenStudy (anonymous):

don't medal until you know XD

OpenStudy (gabbyalicorn):

-_- if i get this wrong.... i dont know what ill do

OpenStudy (anonymous):

unfan?

OpenStudy (anonymous):

idk

OpenStudy (gabbyalicorn):

nah iill just be mad

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

is itttttttt doe

OpenStudy (gabbyalicorn):

okay, um trusting u -.-

OpenStudy (anonymous):

now im stressed ur scaring me

OpenStudy (gabbyalicorn):

ur scaring me XD

OpenStudy (anonymous):

"trust my rage"

OpenStudy (gabbyalicorn):

wait i thought lokis dont get stressed

OpenStudy (anonymous):

shhhh you have a hulk

OpenStudy (gabbyalicorn):

I hate the hulk = least fav marvel character

OpenStudy (anonymous):

fineeeeeee you have a 'venger

OpenStudy (gabbyalicorn):

XD which 1

OpenStudy (anonymous):

natasha?

OpenStudy (gabbyalicorn):

thats not an avenger

OpenStudy (anonymous):

yesssss she is

OpenStudy (anonymous):

widow is a 'venger

OpenStudy (gabbyalicorn):

wait is she the red hair girl?

OpenStudy (anonymous):

duh

OpenStudy (gabbyalicorn):

oh her no1 cares about her

OpenStudy (anonymous):

I DOOOOOOOO AND SO DOES JANEEE

OpenStudy (anonymous):

fine, cap?

OpenStudy (anonymous):

idk u have someone

OpenStudy (gabbyalicorn):

XD yeah, your fine

OpenStudy (gabbyalicorn):

I was replying to , "fine, cap?"

OpenStudy (gabbyalicorn):

yeah, not the other 1

OpenStudy (anonymous):

you got someone so boom, is it right doe? i think it issss.

OpenStudy (gabbyalicorn):

k i trust U

OpenStudy (gabbyalicorn):

i tagged u to the last question come plz